OpenStudy (firejay5):

Simplify. Show work! I have the answer it's hard for me to show my work. 20. y^2 + 4y + 4 over 3y^2 + 5y - 2 21. a^2 + 2a + 1 over 2a^2 + 3a + 1 22. 3xyz over 4xz * 6x^2 over 3y^2 23. -4ab over 21c * 14c^2 over 18a^2

OpenStudy (rc94):

y^2+4y+4 = (y+2)(y+2) 3y^2+5y-2 = (y+2)(3y-1) so you can cancel (y+2) in both to obtain (y+2)/(3y-1)

OpenStudy (rc94):

Is this the sort of work you need, or a little more detail?

OpenStudy (firejay5):

I need the work

OpenStudy (rc94):

I'll try not to give too much away, but in the second example your common factor will be (a+1).

OpenStudy (firejay5):

@RC94 You may give me hints

OpenStudy (rc94):

In the third example make it into one big fraction: 18yzx^3 all over 12xzy^2 Now you can cancel like terms (ie divide top and bottom by 6, and divide top and bottom by xy)

OpenStudy (rc94):

By xyz sorry, not just xy

OpenStudy (rc94):

-56abc^2 all over 378ca^2 Divide top and bottom by 14 and then divide top and bottom by ac and you've got it :)

OpenStudy (firejay5):

@RC94 is #22 3x^2 over 2y for the answer

OpenStudy (rc94):

It is indeed :) seems like you've got the hang of it!

OpenStudy (firejay5):

is the answer to 23. -4c over 27a

OpenStudy (rc94):

Almost it should be -4bc over 27a

OpenStudy (firejay5):

I forgot the b

OpenStudy (rc94):

Yep that's right :)

OpenStudy (firejay5):

add me as a fan, because you are amazing

OpenStudy (firejay5):

How did you get so good